# Defending Against Alignment-Breaking Attacks via Robustly Aligned LLM

### v2

@@ 1 INTRODUCTION
Trained on a wide range of text data from the internet, Large Language Models (LLMs) have exhibited exciting improvement in their generalization capabilities (OpenAI, 2023; Touvron et al., 2023b) and widespread application in various domains such as finance (Wu et al., 2023), law (Nguyen, 2023), and healthcare industry (Thirunavukarasu et al., 2023). While LLMs have showcased impressive potential, a rising concern is that they can also be maliciously utilized to generate content deviating from human values (e.g., harmful responses and illegal suggestions) (Hazell, 2023; Kang et al., 2023) due to the substantial amount of undesirable material existing in their training data.
@@
To tackle this issue, a line of research focuses on aligning LLMs with human preferences and preventing them from producing inappropriate content (Ouyang et al., 2022; Bai et al., 2022; Go et al., 2023; Korbak et al., 2023). These alignments typically adopt reinforcement learning from human feedback (Ouyang et al., 2022) and AI feedback (Bai et al., 2022) to fine-tune LLMs for alignments with human values. Despite these efforts, an emerging class of jailbreak attacks can still bypass the alignment and elicit harmful responses from LLMs (Yuan et al., 2023; Shen et al., 2023; Wei et al., 2023; Zou et al., 2023). These alignment-breaking attacks manually craft adversarial prompts by designing elaborate role-play (Shen et al., 2023) or simply asking the LLM to give the response starting with “Absolutely! Here’s” (Wei et al., 2022). Moreover, automatic jailbreak prompt generation methods have also been developed through dialogue encryption (Yuan et al., 2023) or the combination of greedy and gradient-based search methods (Zou et al., 2023). Figure 1 shows an example that a malicious question appended with an adversarial prompt could successfully break the safety alignment. Recently, (Zou et al., 2023) have demonstrated that jailbreak attempts could be highly effective and transferable across different LLMs. This phenomenon suggests that existing safety alignment is far from robust to defend against carefully crafted adversarial prompts.
Till now, few attempts have been made to design dedicated mechanisms for defending alignment-breaking attacks. A rudimentary defense currently employed relies on external tools to re-assess the potential harm of the LLM responses. For instance, it could feed every potential response from the target LLM into a third-party LLM to determine whether the response is harmful or not (Helbling et al., 2023). While this strategy enables filtering out possible harmful responses, there are several major drawbacks limiting its practicability: 1) Existing LLMs are very sensitive to harmful keywords appeared in the input, and have a high propensity to misclassify benign content as harmful, even when the entire sentence is not talking about any harmful behavior (e.g., stating news or providing guidance/warnings). This could lead to a high false-positive rate in harmful content detection; 2) The method heavily relies on the performance of the LLM used as a harmful discriminator, while the LLM itself is not designed to be an accurate harmful discriminator. The basis for its decisions remains ambiguous, implying that the harmful evaluation process could be opaque; 3) There are more types of alignment that can not be simply summarised as “harmful” (e.g., privacy, ethics, human values etc), thus this type of approach cannot cover such cases simultaneously. Given the wide range of applications where LLMs could be utilized, finding an effective and practical defense against potential alignment-breaking attacks is both urgent and challenging.
@@
Note that our RA-LLM does not require any external “harmful” detectors, instead, our strategy only relies on the existing alignment capability inside the LLM. Due to the same reason, our approach is not limited to any specific type of alignment (e.g., harmful), but robustifies all existing model alignments. Furthermore, we provide a theoretical analysis to verify the effectiveness of our proposed RA-LLM. Our experimental results on open-source large language models demonstrate that RA-LLM can successfully defend against both state-of-the-art adversarial prompts and popular handcrafted jailbreaking prompts by reducing their attack success rates from nearly 100% to around 10% or less.
@@ 2 RELATED WORKS
Foundational large language models are pre-trained on extensive textual corpora (Brown et al., 2020; Chowdhery et al., 2022; Touvron et al., 2023a), which equips LLMs with world knowledge and facilitates their deployment in professional applications. Despite their excellent performance, LLMs suffer from generating outputs that deviate from human expectations (e.g., harmful responses and illegal suggestions) due to the significant amount of inappropriate content existing in unfiltered training data. To tackle this issue, a line of work focuses on aligning LLMs with human values (Xu et al., 2020b; Ouyang et al., 2022; Bai et al., 2022; Go et al., 2023; Korbak et al., 2023). Specifically, Ouyang et al. (2022) align the LLM by using reinforcement learning from human feedback (RLHF (Christiano et al., 2017; Stiennon et al., 2020)) to fine-tune pre-trained LLM with human preferences as the reward signal, which reduces the generation of toxic content. Bai et al. (2022) train a less harmful system through the specification of a short list of principles and further improve the human-judged performance by introducing chain-of-thought style reasoning (Wei et al., 2022) in both supervised-learning and reinforcement-learning stage. Go et al. (2023) consider aligning LLMs as approximating a target distribution representing some desired behavior and accordingly propose a new framework for fine-tuning LLMs to approximate any target distribution through f-divergences minimization. In addition to aligning LLMs in the fine-tuning stage, Korbak et al. (2023) propose pertaining LLMs with alternative objectives that guide them to generate text aligned with human preferences and significantly reduce the rate of generating undesirable content by using conditional training (Keskar et al., 2019).
@@
Although various alignment strategies have been developed to steer LLMs to generate content complying with human ethical principles, an emerging class of alignment-breaking attacks (i.e., jailbreak attacks) can still bypass safeguards and elicit LLMs to generate harmful and toxic responses (Wolf et al., 2023; Li et al., 2023; Shen et al., 2023; Yuan et al., 2023; Wei et al., 2023; Zou et al., 2023), which poses significant threats to the practical deployment of LLMs. In particular, inspired by traditional computer security, Kang et al. (2023) adapt obfuscation, code injection/payload splitting, and visualization attacks to the LLMs, leading to the generation of content containing hate speech, phishing attacks, and scams. Wei et al. (2023) hypothesize that competing objectives and mismatched generalization are two failure modes of safety training in LLMs and craft effective jailbreak attacks by leveraging the two failure modes. Instead of manually crafting adversarial prompts, Zou et al. (2023) automatically produce transferable adversarial suffixes by using greedy and gradient-based search methods to maximize the probability of generating an affirmative response. Yuan et al. (2023) bypass the safety alignment through dialogue encryption. Shen et al. (2023) systematically analyzes the characteristics of jailbreak prompts in the wild and presents that jailbreak prompts have evolved to be more stealthy and effective with reduced length, increased toxicity, and semantic shift. Note that some concurrent works also aim to defend against alignment-breaking attacks: Kumar et al. (2023) provides a verifiable safety guarantee by enumerating all possible partially erased input and using a safety filter to identify the harmfulness of the input content. Jain et al. (2023) propose to detect adversarial prompts by checking if the perplexity of the prompt is greater than a threshold.
@@
Traditional text adversarial attacks primarily focus on text classification tasks and aim to force target models to maximize their prediction error by adversarially perturbing original text (Ebrahimi et al., 2017; Jin et al., 2020; Li et al., 2018; Maheshwary et al., 2021; Ye et al., 2023). The adversarial perturbation could be crafted by performing character-level transformation (Gao et al., 2018) or replacing original words with their synonyms while maintaining semantics and syntax similar (Alzantot et al., 2018). The generation of adversarial examples could be categorized into the “white-box” setting and the “black-box” setting according to the extent of access to the target model (Xu et al., 2020a). As a representative white-box method, HotFlip  (Ebrahimi et al., 2017) uses the gradient information of discrete text structure at its one-hot representation to construct adversarial examples. In the black-box setting, Li et al. (2018); Jin et al. (2020); Ren et al. (2019) leverage the prediction score distribution on all categories to craft adversarial text without the guidance of parameter gradients. Maheshwary et al. (2021) focus on a more realistic scenario where attackers only know the top-$1$ prediction and propose using population-based optimization to construct adversarial text. Ye et al. (2022) follow the same scenario and employ the word embedding space to guide the generation of adversarial examples.
To defend against adversarial attacks, a body of empirical defense methods has been proposed. In particular, adversarial-training-based methods (Miyato et al., 2016; Zhu et al., 2019) incorporate adversarial perturbations to word embeddings and robustly train the model by minimizing the adversarial loss. Zhou et al. (2021); Dong et al. (2021) utilize adversarial data augmentation by replacing the original word with its synonyms to make the model robust to similar adversarial perturbations. These methods gain empirical success against adversarial attacks. To provide provable robustness against adversarial word substitutions, Jia et al. (2019) use certifiably robust training by training the model to optimize Interval Bound Propagation (IBP) upper bound. Shi et al. (2020) adopt linear-relaxation-based perturbation analysis (Xu et al., 2020c) to develop a robustness verification method for transformers. Zeng et al. (2023) propose a certifiably robust defense method based on randomized smoothing techniques (Cohen et al., 2019).
@@ 3.1 Threat Model
We also assume that the target LLM $f(\cdot)$ is an already aligned LLM that has a certain ability to reject commonly seen malicious requests. In other words, when the malicious question ${\mathbf{x}}$ is directly input into the target LLM $f(\cdot)$, it will, in most cases, deny answering such a question by outputting a response similar to “I am sorry, but I cannot talk about [a malicious request]…”. On the contrary, the alignment-breaking attacker’s goal is to break the existing alignment of the target LLM by finding an adversarial prompt ${\mathbf{p}}_{\text{adv}}$, so that ${\mathbf{x}}_{\text{adv}}={\mathbf{x}}\oplus{\mathbf{p}}_{\text{adv}}$ will mislead the LLM to provide an affirmative answer Zou et al. (2023) to such a malicious question, e.g., “Sure, here is how to do [a malicious request]…”.
@@ 3.2 Our Proposed Method
$f^{\prime}({\mathbf{x}})=\left\{\begin{aligned} &\text{Reject the response},\text{ if }\text{AC}(f({\mathbf{x}}))=\text{Fail}\\ &f({\mathbf{x}})\ \ \ \ \ \ \ \qquad\qquad,\text{ if }\text{AC}(f({\mathbf{x}}))=\text{Pass}\end{aligned}\right.$ (1)
@@
One thing to notice here is that the previously defined alignment check function $\text{AC}(\cdot)$ only relies on the existing alignments inside on the target LLM. However, the existence of alignment-breaking attacks such as the adversarial prompts Zou et al. (2023) has proved that such alignment checking is not robust: it can be easily manipulated and circumvented by carefully designed perturbations or suffix prompts. Therefore, it is natural to think about how we can design a robust alignment check function that could strengthen the alignment check capabilities of an aligned LLM, without finetuning or modifying the model itself.
@@
Eq. 2 states that the robust alignment check function $\text{RAC}(\cdot)$ not only requires the original response $f({\mathbf{x}})$ to show no sign of being aligned but also requires the response after random dropping still shows no sign of being aligned in most cases. On the contrary, if $\text{AC}({\mathbf{x}})$ already fails or over a certain ratio (e.g., $>t$) of responses from the randomly dropped input fails to pass AC, $\text{RAC}(\cdot)$ will also fail it. Therefore, it is easy to see that such a design certainly helps us build a more robust alignment check function compared to $\text{AC}(\cdot)$.
@@
$f_{\text{rob}}({\mathbf{x}})=\left\{\begin{aligned} &\text{Reject the response},\text{ if }\text{RAC}(f({\mathbf{x}}))=\text{Fail}\\ &f({\mathbf{x}})\ \ \ \ \ \ \ \qquad\qquad,\text{ if }\text{RAC}(f({\mathbf{x}}))=\text{Pass}\end{aligned}\right.$ (3)
@@ 3.3 Practical Designs
Previously, we vaguely defined the alignment check function $\text{AC}(\cdot)$ as returning Fail when detecting typical aligned output while returning Pass otherwise. In practice, we approximate this alignment check function through prefix checking: we observed that various aligned outputs often share similar prefixes such as “I can not”, “I’m sorry”. Therefore, we can build a prefix set and if any prefix in the set appears in LLM’s output, the alignment check function $\text{AC}(\cdot)$ returns Fail; otherwise, it returns Pass. Note that we are only inspecting the prefix. For this purpose, we only need to generate a certain number of tokens (e.g., 10) for robust alignment checking. This could largely reduce our computational overhead [^2] Further discussion on computational costs can be found in Section 4.5..
@@ 3.4 Theoretical Analysis
In this section, we theoretically analyze the proposed robustly aligned LLM and see when it provides a more robust alignment compared to the original LLM when facing alignment-breaking attacks.
Our theorem is based on the analysis on the robust alignment check function RAC. We will show that RAC is more robust for the aligned malicious text $x$ with any adversarial prompt ${\mathbf{p}}_{\text{adv}}$ of length $M$ and it can be inserted into any position (e.g., in front, back, or middle of ${\mathbf{x}}$).
@@
[…]

### v3

@@ Abstract
Defending Against Alignment-Breaking Attacks via Robustly Aligned LLM
Bochuan Caothanks: Equal Contribution , Yuanpu Cao11footnotemark: 1 , Lu Lin & Jinghui Chen The Pennsylvania State University {bccao,ymc5533,lulin,jzc5917}@psu.edu
@@ 1 INTRODUCTION
Trained on a wide range of text data from the internet, Large Language Models (LLMs) have exhibited exciting improvement in their generalization capabilities (OpenAI, 2023; Touvron et al., 2023) and widespread application in various domains such as finance (Wu et al., 2023), law (Nguyen, 2023), and healthcare industry (Thirunavukarasu et al., 2023). While LLMs have showcased impressive potential, a rising concern is that they can also be maliciously utilized to generate content deviating from human values (e.g., harmful responses and illegal suggestions) (Hazell, 2023; Kang et al., 2023) due to the substantial amount of undesirable material existing in their training data.
@@
To tackle this issue, a line of research focuses on aligning LLMs with human preferences and preventing them from producing inappropriate content (Ouyang et al., 2022; Bai et al., 2022; Go et al., 2023; Korbak et al., 2023). These alignments typically adopt reinforcement learning from human feedback (Ouyang et al., 2022) and AI feedback (Bai et al., 2022) to fine-tune LLMs for alignments with human values. Despite these efforts, an emerging class of jailbreak attacks can still bypass the alignment and elicit harmful responses from LLMs (Yuan et al., 2023; Shen et al., 2023; Wei et al., 2023a; Zou et al., 2023). These alignment-breaking attacks manually craft adversarial prompts by designing elaborate role-play (Shen et al., 2023) or simply asking the LLM to give the response starting with “Absolutely! Here’s" (Wei et al., 2022). Moreover, automatic jailbreak prompt generation methods have also been developed through dialogue encryption (Yuan et al., 2023) or the combination of greedy and gradient-based search methods (Zou et al., 2023). Figure 1 shows an example that a malicious question appended with an adversarial prompt could successfully break the safety alignment. Recently, Zou et al. (2023) have demonstrated that jailbreak attempts could be highly effective and transferable across different LLMs. This phenomenon suggests that existing safety alignment is far from robust to defend against carefully crafted adversarial prompts.
Till now, few attempts have been made to design dedicated mechanisms for defending alignment-breaking attacks. A rudimentary defense currently employed relies on external tools to re-assess the potential harm of the LLM responses. For instance, it could feed every potential response from the target LLM into a third-party LLM to determine whether the response is harmful or not (Helbling et al., 2023). While this strategy enables filtering out possible harmful responses, there are several major drawbacks limiting its practicability: 1) Existing LLMs are very sensitive to harmful keywords appeared in the input, and have a high propensity to misclassify benign content as harmful, even when the entire sentence is not talking about any harmful behavior (e.g., stating news or providing guidance/warnings). This could lead to a high false-positive rate in harmful content detection; 2) The method heavily relies on the performance of the LLM used as a harmful discriminator, while the LLM itself is not designed to be an accurate harmful discriminator. The basis for its decisions remains ambiguous, implying that the harmful evaluation process could be opaque; 3) Some types of alignment that can not be simply summarised as “harmful” (e.g., privacy, ethics, human values etc), thus it cannot cover such cases simultaneously. Given the wide range of applications where LLMs could be utilized, finding an effective and practical defense against potential alignment-breaking attacks is both urgent and challenging.
@@
Note that our RA-LLM does not require any external “harmful” detectors, instead, our strategy only relies on the existing alignment capability inside the LLM. Due to the same reason, our approach is not limited to any specific type of alignment (e.g., harmful), but robustifies all existing model alignments. Furthermore, we provide a theoretical analysis to verify the effectiveness of RA-LLM. Our experimental results on open-source large language models demonstrate that RA-LLM can successfully defend against both state-of-the-art adversarial prompts and popular handcrafted jailbreaking prompts by reducing their attack success rates from nearly 100% to around 10% or less.
@@ 2 RELATED WORKS
Despite the excellent generalization capabilities, LLMs suffer from generating outputs that deviate from human expectations due to the significant amount of inappropriate content existing in unfiltered training data. To tackle this issue, a line of work focuses on aligning LLMs with human values (Xu et al., 2020b; Ouyang et al., 2022; Bai et al., 2022; Go et al., 2023; Korbak et al., 2023). Specifically, Ouyang et al. (2022) align LLMs by using reinforcement learning from human feedback (Christiano et al., 2017; Stiennon et al., 2020) to fine-tune pre-trained LLM with human preferences as the reward signal, which reduces the generation of toxic content. Bai et al. (2022) train a less harmful system through the specification of a short list of principles and further improve the human-judged performance by introducing chain-of-thought style reasoning (Wei et al., 2022) in fine-tuning stage. In addition to aligning LLMs in the fine-tuning stage, Korbak et al. (2023) propose pertaining LLMs with alternative objectives that guide them to generate text aligned with human preferences and significantly reduce the generation of undesirable content by using conditional training (Keskar et al., 2019).
@@
Although various alignment strategies have been developed to steer LLMs to generate content complying with human ethical principles, an emerging class of alignment-breaking attacks (i.e., jailbreak attacks) can still bypass safeguards and elicit LLMs to generate harmful and toxic responses (Li et al., 2023; Shen et al., 2023; Yuan et al., 2023; Cao et al., 2023; Kang et al., 2023; Zou et al., 2023; Wei et al., 2023b; Helbling et al., 2023), which poses significant threats to the practical deployment of LLMs. In particular, inspired by traditional computer security, Kang et al. (2023) adapt obfuscation, code injection/payload splitting, and visualization attacks to LLMs, leading to the generation of content containing hate speech, phishing attacks, and scams. Instead of manually crafting adversarial prompts, Zou et al. (2023) automatically produce transferable adversarial suffixes by using greedy and gradient-based search to maximize the probability of generating an affirmative response. AutoDAN (Liu et al., 2023) also can automatically generate jailbreak prompts through a genetic algorithm with a handcrafted jailbreak prompt as initialization. TAP (Mehrotra et al., 2023) iteratively refine the candidate attack prompts using tree-of-thought reasoning.
Note that some concurrent works also aim to defend against alignment-breaking attacks: Kumar et al. (2023) provides a verifiable safety guarantee by enumerating all possible partially erased input and using a safety filter to identify the harmfulness of the input content. LLM self-defense (Helbling et al., 2023) simply utilizes itself or another LLM to detect if its own response is harmful. Jain et al. (2023) propose to detect adversarial prompts by checking if the perplexity of the prompt is greater than a threshold. We defer more discussion and comparison with concurrent defense methods in Appendix F and G.
@@
Traditional text adversarial attacks primarily focus on text classification tasks and aim to force target models to maximize their prediction error by adversarially perturbing original text (Ebrahimi et al., 2017; Jin et al., 2020; Li et al., 2018; Maheshwary et al., 2021; Ye et al., 2023). The adversarial perturbation could be crafted by performing character-level transformation (Gao et al., 2018) or replacing original words with their synonyms while maintaining semantics similar (Alzantot et al., 2018). The generation of adversarial examples could be categorized into the “white-box" setting and the “black-box" setting according to the extent of access to the target model (Xu et al., 2020a). As a representative white-box method, HotFlip  (Ebrahimi et al., 2017) uses the gradient information of discrete text structure at its one-hot representation to construct adversarial examples. In the black-box setting, Li et al. (2018); Jin et al. (2020); Ren et al. (2019) leverage the prediction score distribution on all categories to craft adversarial text without the guidance of parameter gradients. Maheshwary et al. (2021) focus on a more realistic scenario where attackers only know the top-$1$ prediction and propose using population-based optimization to construct adversarial text.
To defend against adversarial attacks, a body of empirical defense methods has been proposed. In particular, adversarial-training-based methods (Miyato et al., 2016; Zhu et al., 2019) incorporate adversarial perturbations to word embeddings and robustly train the model by minimizing the adversarial loss. Zhou et al. (2021); Dong et al. (2021) utilize adversarial data augmentation by replacing the original word with its synonyms to make the model robust to similar adversarial perturbations. To provide provable robustness against adversarial word substitutions, Jia et al. (2019) use certifiably robust training by training the model to optimize Interval Bound Propagation (IBP) upper bound. Shi et al. (2020) adopt linear-relaxation-based perturbation analysis (Xu et al., 2020c) to develop a robustness verification method for transformers. Zeng et al. (2023) propose a certifiably robust defense method based on randomized smoothing techniques (Cohen et al., 2019).
@@ 3.1 Threat Model
We also assume that the target LLM $f(\cdot)$ is an already aligned LLM that has a certain ability to reject commonly seen malicious requests. In other words, when the malicious question ${\mathbf{x}}$ is directly input into the target LLM $f(\cdot)$, it will, in most cases, deny answering such a question by outputting a response similar to “I am sorry, but I cannot talk about [a malicious request]…”. On the contrary, the attacker’s goal is to break the existing alignment of the target LLM by finding an adversarial prompt ${\mathbf{p}}_{\text{adv}}$, so that ${\mathbf{x}}_{\text{adv}}={\mathbf{x}}\oplus{\mathbf{p}}_{\text{adv}}$ will mislead the LLM to provide an affirmative answer Zou et al. (2023) to such a malicious question, e.g., “Sure, here is how to do [a malicious request]…”.
@@ 3.2 Our Proposed Method
$f^{\prime}({\mathbf{x}})=\left\{\begin{aligned} &\text{Reject to answer},\text{ if }\text{AC}(f({\mathbf{x}}))=\text{Fail}\\ &f({\mathbf{x}})\ \ \ \ \ \ \ \quad\qquad,\text{ if }\text{AC}(f({\mathbf{x}}))=\text{Pass}\end{aligned}\right.$ (1)
@@
One thing to notice here is that the previously defined alignment check function $\text{AC}(\cdot)$ only relies on the existing alignments inside of the target LLM. However, the existence of alignment-breaking attacks such as the adversarial prompts Zou et al. (2023) has proved that such alignment checking is not robust: it can be easily manipulated and circumvented by carefully designed perturbations or suffix prompts. Therefore, it is natural to think about how we can design a robust alignment check function that could strengthen the alignment check capabilities of an aligned LLM, without finetuning or modifying the model itself.
@@
Eq. 2 states that the robust alignment check function $\text{RAC}(\cdot)$ not only requires the original response $f({\mathbf{x}})$ to show no sign of being aligned (e.g. refusal-to-answer) but also requires the response after random dropping still shows no sign of being aligned in most cases. Typically, aligned text for malicious requests includes some refusal-to-answer text, such as "I am sorry, but I cannot answer this…". On the contrary, if $\text{AC}({\mathbf{x}})$ already fails or over a certain ratio (e.g., $>t$) of responses from the randomly dropped input fails to pass AC, $\text{RAC}(\cdot)$ will also fail it. Therefore, it is easy to see that such a design certainly helps us build a more robust alignment check function compared to $\text{AC}(\cdot)$.
@@
$f_{\text{rob}}({\mathbf{x}})=\left\{\begin{aligned} &\text{Reject to answer},\text{if }{\text{RAC}(f({\mathbf{x}}))=\text{Fail}}\\ &f({\mathbf{x}})\ \ \ \ \ \ \quad\qquad,\text{ if }{\text{RAC}(f({\mathbf{x}}))=\text{Pass}}\end{aligned}\right.$ (3)
@@ 3.3 Practical Designs
Previously, we vaguely defined the alignment check function $\text{AC}(\cdot)$ as returning Fail when detecting typical aligned output while returning Pass otherwise. In practice, we approximate this alignment check function through prefix checking: we observed that various aligned outputs often share similar prefixes such as “I can not”, “I’m sorry”. Therefore, we can build a prefix set and if any prefix in the set appears in LLM’s output, the alignment check function $\text{AC}(\cdot)$ returns Fail; otherwise, it returns Pass. Note that we are only inspecting the prefix. For this purpose, we only need to generate a certain number of tokens (e.g., 10) for robust alignment checking. This could largely reduce our computational overhead [^2] Further discussion on computational costs can be found in Section H.2..
@@ 3.4 Theoretical Analysis
In this section, we theoretically analyze the proposed robustly aligned LLM and see when it provides a more robust alignment compared to the original LLM when facing alignment-breaking attacks. Our theorem is based on the analysis on the robust alignment check function RAC. We will show that RAC is more robust for the aligned malicious text $x$ with any adversarial prompt ${\mathbf{p}}_{\text{adv}}$ of length $M$ and it can be inserted into any position (e.g., in front, back, or middle of ${\mathbf{x}}$).
@@
Consider a malicious input ${\mathbf{x}}$ and its corresponding adversarial prompt ${\mathbf{p}}_{\text{adv}}$ such that ${\mathbf{x}}_{\text{adv}}={\mathbf{x}}\oplus{\mathbf{p}}_{\text{adv}}$ could break the alignment in the LLM $f(\cdot)$. Suppose ${\mathbf{x}}$ consists of $N$ tokens and ${\mathbf{p}}_{\text{adv}}$ consists of $M$ tokens while ${\mathbf{p}}_{\text{adv}}$ could be insert to any position $j\in[0,...,N]$ in ${\mathbf{x}}$. Denote ${\mathbf{x}}_{\text{pad}}^{j}$ as the padded text constructed from ${\mathbf{x}}$ by inserting $M$ pad tokens into position $j$. If $N\geq\frac{M(1-p)}{p}$ and $\mathop{\min}\limits_{j}\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U(p)}(\text{AC}(f([{\mathbf{x}}_{\text{pad}}^{j}]_{{\mathbf{r}}}))=\text{Fail})>t+c,$ where $c=1-\frac{\binom{N}{(N+M)(1-p)}}{\binom{N+M}{(N+M)(1-p)}}$ and $t$ is the threshold used in Algorithm 1, then our robustly aligned LLM in Algorithm 1 with sufficiently large random drop trials $n$ will reject the request on ${\mathbf{x}}_{\text{adv}}={\mathbf{x}}\oplus{\mathbf{p}}_{\text{adv}}$.